# Space spanned by characteristic exponents

### v2

@@ Abstract.
We prove several rigidity results on multiplier spectrum and length spectrum. For example, we show that for every non-exceptional rational map $f:\mathbb{P}^{1}(\mathbb{C})\to\mathbb{P}^{1}(\mathbb{C})$ of degree $d\geq 2$, the $\mathbb{Q}$-vector space generated by all the (finite) characteristic exponents of periodic points of $f$ has infinite dimension. This answers a stronger version of a question of Levy and Tucker. Our result can also be seen as a generalization of recent results of Ji-Xie and of Huguin which proved Milnor’s conjecture about rational maps having integer multipliers. We also get a characterization of postcritically finite maps by using its length spectrum. Finally as an application of our result, we get a new proof of the Zariski-dense orbit conjecture for endomorphisms on $(\mathbb{P}^{1})^{N},N\geq 1$.
@@ 1.1. Multiplier, length, and characteristic exponent
Let $z_{0}\in\mathbb{P}^{1}(\mathbb{C})$ be a periodic point of $f$ with exact period $n$. Define $n_{f}(z_{0}):=n$ be this period. We write $n(z_{0})$ for simplicity when the map $f$ is clear. The multiplier $\rho_{f}(z_{0})$ of $f$ at $z_{0}$ is defined to be the differential $df^{n}(z_{0})\in\mathbb{C}$. We write $\rho(z_{0})$ for simplicity when the map $f$ is clear. The length of $f$ at $z_{0}$ is the norm $|\rho_{f}(z_{0})|$. The multiplier and the length are invariant under conjugacy. The characteristic exponent of $f$ at $z_{0}$ is defined to be $\chi_{f}(z_{0}):=n^{-1}\log\lvert\rho_{f}(z_{0})\rvert$ (when $\rho_{f}(z_{0})\neq 0$).
@@
where $\mu_{f}$ is the unique maximal entropy measure of $f$ on $\mathbb{P}^{1}(\mathbb{C})$ which is also the canonical measure of $f$ on $\mathbb{P}^{1}(\mathbb{C})$ (see [14, 7, 15, 34]), and the norm of the differential is computed with respect to the spherical metric. In fact, the Lyapunov exponent dose not change with the spherical metric replaced by any Kähler metric on $\mathbb{P}^{1}_{\mathbb{C}}$ in the definition. It is known that $\mathcal{L}_{f}\geq\log(d)/2$, and the equality holds if and only if $f$ is a Lattès map in the sense of Definition 1.1 (see [34, 35]). The quantity $\exp(\mathcal{L}_{f})$ is considered as an average of $|df|$ on $\mathbb{P}^{1}(\mathbb{C})$, which measures the average rate of expansion of $f$ along a typical orbit.
@@ 1.2. Exceptional maps
  • The map $f$ is called Lattès if there is an endomorphism $\phi$ on an elliptic curve $E$ such that $\phi$ is semi-conjugate to $f$. Here, for an endomorphism $g$ on an algebraic variety $X$ and an endomorphism $h$ on an algebraic variety $Y$, we say $g$ is semi-conjugate to $h$ if there is a dominant morphism $\pi:X\to Y$ such that $\pi\circ g=h\circ\pi$. Furthermore, the map $f$ is called flexible Lattès if there are an elliptic curve $E$ and $n\in\mathbb{Z}\setminus\{0,\pm 1\}$ such that the multiplication $[n]$ on $E$ is semi-conjugate to $f$ via the quotient map modulo {±$\{\pm 1\}$ on $E$. Otherwise, it is called rigid Lattès.
  • We say that $f$ is of monomial type if it semi-conjugate to the map $z\mapsto z^{n}$ on $\mathbb{P}^{1}$ for some integer $n\neq 0,\pm 1$.
@@ 1.3. Statement of the main results
We fix an embedding of the algebraic closure $\overline{\mathbb{Q}}$ of $\mathbb{Q}$ in $\mathbb{C}$ and consider all number fields as subfields of $\mathbb{C}$. Denote the usual absolute value on $\mathbb{C}$ by ||$\cdot$|| ⋅ |.
@@
In Theorem 2.1, we prove a more general version of Theorem 1.2, in which $\overline{\mathbb{Q}}$ can be replaced by any algebraically closed subfield of $\mathbb{C}$ invariant under the complex conjugation. In fact, the actual proof shows that we can weaken the conditions in Theorems 1.2 and 2.1 by discarding the corresponding condition for finitely many periodic points of $f$, see Remark 2.2. For example, in Theorem 1.2, we can only require that $|\rho_{f}(x)|\in\overline{\mathbb{Q}}$ for all but finitely many $x\in{\rm Per}(f)(\mathbb{C})$, and the conclusion still holds.
McMullen’s rigidity of multiplier spectrum [16, Corollary 2.3] asserts that multipliers at all periodic points determine the conjugacy class of rational maps up to only finitely many choices, except for the flexible Lattès family. With a standard spread out argument, McMullen’s rigidity theorem implies that, for a rational map $f$ of degree at least $2$ which is not flexible Lattès, if its multipliers at periodic points are all algebraic, then $f$ is defined over $\overline{\mathbb{Q}}$. We sketch the argument as follows. Let $f$ be a rational map over $\mathbb{C}$ of degree $d\geq 2$ that is not flexible Lattès with all multipliers in $\overline{\mathbb{Q}}$. The noetherianity implies that the set $D$ of conjugate classes $[g]$ (in the moduli space $\mathcal{M}_{d}$) such that $g$ has the same multiplier spectrum (i.e., all multipliers at periodic points) as $f$ is an algebraic subset of $\mathcal{M}_{d}$ defined over $\overline{\mathbb{Q}}$, since the multiplier spectrum morphism is defined over $\mathbb{Q}$ and all multipliers of $f$ are in $\overline{\mathbb{Q}}$. Then McMullen’s rigidity implies that $\dim(D)=0$, hence $D$ is a set of finitely many $\overline{\mathbb{Q}}$-points. Thus $f$ is defined over $\overline{\mathbb{Q}}$ as $[f]\in D$.
Theorem 1.2 is a generalization of this result from multiplier spectrum to length spectrum (which contains less information). The rigidity of length spectrum was proved by Ji-Xie [12, Theorem 1.5]. However, the spread out argument does not apply directly in this case as the length spectrum map (and its square) is not algebraic on the moduli space of rational maps. Indeed as shown in [12, §8.1], its square is not even real algebraic (but it is still semialgebraic). In §2.3, we introduce a method to do the spread out argument respecting the structure of $\mathcal{M}_{d}(\mathbb{C})$ as a real algebraic variety using Weil restriction. Another difficulty in the length spectrum case is the lack of noetherianity for semialgebraic subsets. We overcome this difficulty using the notion of admissible subsets introduced in [12].
@@
With the help of Remark 1.3, the proof of Theorem 1.5 in §4 indeed shows that we can weaken the condition of Theorem 1.5 by only requiring that there is a number field $K$ such that for all but finitely many $z_{0}\in{\rm Per}(f)$, we have $\lvert\rho_{f}(z_{0})\rvert^{n(z_{0})}\in K$ for some $n(z_{0})\in\mathbb{Z}_{>0}$.
Finitely many nonzero elements $z_{1},\dots,z_{N}$ in a field $F$ are called multiplicatively independent if for all triples $(m_{1},\dots,m_{N})$ of integers, $z_{1}^{m_{1}}\cdots z_{N}^{m_{N}}=1$ if and only if $m_{1}=\cdots=m_{N}=0$. A sequence $(z_{n})_{n=1}^{\infty}$ in $F\setminus\{0\}$ is called multiplicatively independent if all its finite subsequences are multiplicatively independent. Theorem 1.4 immediately implies the existence of infinitely many multipliers for a non-exceptional $f$ whose absolute values are multiplicatively independent.
@@ 1.4.1. Milnor’s conjecture
Milnor [18] has showed that an exceptional rational map $f:\mathbb{P}^{1}(\mathbb{C})\to\mathbb{P}^{1}(\mathbb{C})$ of degree $d\geq 2$ must have all its multipliers of periodic points in the ring of integers $\mathcal{O}_{K}$ for some imaginary quadratic number field $K$, and in fact in $\mathbb{Z}$ when $f$ is not a rigid Lattès map. Milnor conjectured that the converse is also true. Milnor’s conjecture was recently proved by Ji-Xie:
@@
See also [5] for a different proof. Recently Huguin generalized the above result using different approach:
@@ 1.4.2. A question of Levy and Tucker
On the other hand, in the 2014 AIM workshop Postcritically Finite Maps In Complex And Arithmetic Dynamics, Levy [13] and Tucker [26] asked the following question independently:
@@
Let $f:\mathbb{P}^{1}(\mathbb{C})\to\mathbb{P}^{1}(\mathbb{C})$ be a non-exceptional rational map of degree $d\geq 2$ and let $S$ be the set of all multipliers of periodic points of $f$. Take the subgroup of $\mathbb{C}^{\ast}$ generated by $S\setminus\{0\}$. Is that true that the rank of this group is infinite?
@@ 1.5. Sketch of the proofs
We have given the idea of Theorem 1.2 before. Here we explain the proofs of Theorem 1.4 and Theorem 1.5.
We first give the idea of the proof of Theorem 1.5. We argue by contradiction and suppose that $f$ is not exceptional. The first step is to reduce to the case where $f$ is defined over $\overline{\mathbb{Q}}$, which can be done using our Theorem 1.2. After enlarging $K$, we may assume that both $f$ and $\overline{f}$ are defined over $K$, where $\overline{f}$ is the rational map obtained from $f$ via replacing the coefficients by their complex conjugations. In the second step, we combine an arithmetic equidistribution theorem with a result of Zdunik [35] (see [10]) on the Lyapunov exponent to get a contradiction. This argument is inspired by Huguin’s proof of Theorem 1.9. Contrary to the case of Theorem 1.9, we can not apply the equidistribution theorem to the one-dimensional dynamical system $f:\mathbb{P}^{1}\to\mathbb{P}^{1}$ directly. Our idea is to consider the two-dimensional system $F:=f\times\overline{f}$ on $\mathbb{P}^{1}\times\mathbb{P}^{1}$ instead. More precisely, applying a result of Zdunik [35], we get a sequence $(x_{n})_{n=1}^{\infty}$ of distinct periodic points of $f$ such that
@@
Consider the endomorphism $F:=f\times\overline{f}$ on $\mathbb{P}^{1}\times\mathbb{P}^{1}$ and $\Gamma:=\overline{\{p_{n}=(x_{n},\overline{x_{n}})\}}^{\rm Zar}\subseteq\mathbb{P}^{1}\times\mathbb{P}^{1}$. Without loss of generality, we may assume that $\Gamma$ is irreducible and that the sequence $(p_{n})_{n=1}^{\infty}$ is generic in $\Gamma$. Since the Dynamical Manin-Mumford problem is solved for such an endomorphism $F$ by [8] and $\Gamma$ contains a Zariski-dense subset $\{p_{n}:n\geq 1\}$ of $F$-periodic points, the subscheme $\Gamma$ is $F$-periodic itself. Hence we may assume that $\Gamma$ is $F$-invariant after replacing $F$ by a suitable iteration.
Let $\nu_{n}$ be the discrete probability measure equally supported at the union of Galois orbits of iterates of $p_{n}$ under $F$. Then $\nu_{n}$ converges weakly to the canonical measure $\mu$ on $\Gamma$ with respect to $F$ by an equidistribution-type theorem (Theorem 3.1), which is a reformulation of [32, Theorem 3.1], see §3 for details. Applying $\nu_{n}\to\mu$ to the continuous test functions $\max\{\log\lvert\det(dF)\rvert,A\}$ ($A\in\mathbb{R}$) and letting $A\to-\infty$, we get
[…]
which is impossible since the right hand side equals to $2\mathcal{L}_{f}<2a$ by a direct computation.
Next we sketch the proof of Theorem 1.4. According to [6], postcritically finite (PCF) maps are defined over $\overline{\mathbb{Q}}$ in the moduli space $\mathcal{M}_{d}$ of rational maps of degree $d$, except for the family of flexible Lattès maps. So it suffices to consider the following two cases: 1). $f$ is defined over $\overline{\mathbb{Q}}$, and 2). $f$ is not PCF. For the first case the conclusion follows from Theorem 1.5. For the second case, we need to develop some new techniques, which are presented in §5. In §5, we consider some pseudo-linear algebra (which means that the domain of a map may not be the whole vector space), and the vector space $\mathbb{D}(k)_{\mathbb{Q}}=k^{*}\otimes_{\mathbb{Z}}\mathbb{Q}$ for a field $k$ of characteristic zero. We will actually prove a theorem (Theorem 5.7) stronger than the non-PCF case of Theorem 1.4, see §5 and §6 for details. To prove Theorem 5.7, in §6.1 we first deal with the case that $f$ is defined over $\overline{\mathbb{Q}}$. A key ingredient in this step is [3, Lemma 4.1] which is a consequence of Siegel’s theorem on $S$-integral points. The existence of a non-preperiodic critical point of $f$ is essentially used here. In §6.2, we consider the general case and finish the proof. This is achieved by reducing to the case that $f$ is defined over $\overline{\mathbb{Q}}$ via an algebraic-geometric argument and techniques developed in §5.

### v3

@@ Abstract.
We prove several rigidity results on multiplier spectrum and length spectrum. For example, we show that for every non-exceptional rational map $f:\mathbb{P}^{1}(\mathbb{C})\to\mathbb{P}^{1}(\mathbb{C})$ of degree $d\geq 2$, the $\mathbb{Q}$-vector space generated by all the (finite) characteristic exponents of periodic points of $f$ has infinite dimension. This answers a stronger version of a question by Levy and Tucker. Our result can also be seen as a generalization of recent results of Ji-Xie and of Huguin which proved Milnor’s conjecture about rational maps having integer multipliers. We also get a characterization of postcritically finite maps by using their length spectra. Finally as an application of our result, we get a new proof of the Zariski-dense orbit conjecture for endomorphisms on $(\mathbb{P}^{1})^{N},N\geq 1$.
@@ 1.1. Multiplier, length, and characteristic exponent
Let $z_{0}\in\mathbb{P}^{1}(\mathbb{C})$ be a periodic point of $f$ with exact period $n$. Define $n_{f}(z_{0}):=n$. We write $n(z_{0})$ for simplicity when the map $f$ is clear. The multiplier $\rho_{f}(z_{0})$ of $f$ at $z_{0}$ is defined to be the differential $df^{n}(z_{0})\in\mathbb{C}$. We write $\rho(z_{0})$ for simplicity when the map $f$ is clear. The length of $f$ at $z_{0}$ is the absolute value $|\rho_{f}(z_{0})|$. The multiplier and length are invariant under conjugacy. The characteristic exponent of $f$ at $z_{0}$ is defined to be $\chi_{f}(z_{0}):=n^{-1}\log\lvert\rho_{f}(z_{0})\rvert$ (when $\rho_{f}(z_{0})\neq 0$).
@@
where $\mu_{f}$ is the unique maximal entropy measure of $f$ on $\mathbb{P}^{1}(\mathbb{C})$ which is also the canonical measure of $f$ on $\mathbb{P}^{1}(\mathbb{C})$ (see [14, 7, 15, 34]), and the norm of the differential is computed with respect to the spherical metric. In fact, the value of the Lyapunov exponent does not change with the spherical metric replaced by any Kähler metric on $\mathbb{P}^{1}_{\mathbb{C}}$ in the definition. It is known that $\mathcal{L}_{f}\geq\log(d)/2$, and the equality holds if and only if $f$ is a Lattès map in the sense of Definition 1.1 (see [34, 35]). The quantity $\exp(\mathcal{L}_{f})$ is considered as an average of $|df|$ on $\mathbb{P}^{1}(\mathbb{C})$, which measures the average expansion rate of $f$ along a typical orbit.
@@ 1.2. Exceptional maps
  • The map $f$ is called Lattès if there is an endomorphism $\phi$ on an elliptic curve $E$ such that $\phi$ is semi-conjugate to $f$. Here, for an endomorphism $g$ on an algebraic variety $X$ and an endomorphism $h$ on an algebraic variety $Y$, we say $h$ is semi-conjugate to $g$ if there is a dominant morphism $\pi:X\to Y$ such that $\pi\circ g=h\circ\pi$. Furthermore, the map $f$ is called flexible Lattès if there are an elliptic curve $E$ and $n\in\mathbb{Z}\setminus\{0,\pm 1\}$ such that $f$ is semi-conjugate to the multiplication-by-$n$ map $[n]$ on $E$ via the quotient map modulo $\{\pm 1\}$ on $E$. Otherwise, it is called rigid Lattès.
  • We say that $f$ is of monomial type if it is semi-conjugate to the map $z\mapsto z^{n}$ on $\mathbb{P}^{1}$ for some integer $n\neq 0,\pm 1$.
@@ 1.3. Statement of the main results
We fix an embedding of the algebraic closure $\overline{\mathbb{Q}}$ of $\mathbb{Q}$ in $\mathbb{C}$ and consider all number fields as subfields of $\mathbb{C}$. Denote the usual absolute value on $\mathbb{C}$ by $|\cdot|$.
@@
In Theorem 2.1, we prove a more general version of Theorem 1.2, in which $\overline{\mathbb{Q}}$ can be replaced by any algebraically closed subfield of $\mathbb{C}$ invariant under complex conjugation. In fact, the actual proof shows that we can weaken the conditions in Theorems 1.2 and 2.1 by discarding the corresponding condition for finitely many periodic points of $f$; see Remark 2.2. For example, in Theorem 1.2, we can require only that $|\rho_{f}(x)|\in\overline{\mathbb{Q}}$ for all but finitely many $x\in{\rm Per}(f)(\mathbb{C})$, and the conclusion still holds.
McMullen’s rigidity of multiplier spectrum [16, Corollary 2.3] asserts that multipliers at all periodic points determine the conjugacy class of rational maps up to only finitely many choices, except for the flexible Lattès family. With a standard spread-out argument, McMullen’s rigidity theorem implies that, for a rational map $f$ of degree at least $2$ which is not flexible Lattès, if its multipliers at periodic points are all algebraic, then $f$ is defined over $\overline{\mathbb{Q}}$. We sketch the argument as follows. Let $f$ be a rational map over $\mathbb{C}$ of degree $d\geq 2$ that is not flexible Lattès with all multipliers in $\overline{\mathbb{Q}}$. Noetherianity implies that the set $D$ of conjugacy classes $[g]$ (in the moduli space $\mathcal{M}_{d}$) such that $g$ has the same multiplier spectrum (i.e., all multipliers at periodic points) as $f$ is an algebraic subset of $\mathcal{M}_{d}$ defined over $\overline{\mathbb{Q}}$, since the multiplier spectrum morphism is defined over $\mathbb{Q}$ and all multipliers of $f$ are in $\overline{\mathbb{Q}}$. Then McMullen’s rigidity implies that $\dim(D)=0$, hence $D$ is a set of finitely many $\overline{\mathbb{Q}}$-points. Thus $f$ is defined over $\overline{\mathbb{Q}}$ as $f]\in D$.
Theorem 1.2 is a generalization of this result from multiplier spectrum to length spectrum (which contains less information). The rigidity of length spectrum was proved by Ji-Xie [12, Theorem 1.5]. However, the spread-out argument does not apply directly in this case as the length spectrum map (and its square) is not algebraic on the moduli space of rational maps. Indeed, as shown in [12, §8.1], its square is not even real algebraic (though it is still semialgebraic). In §2.3, we introduce a method to do the spread-out argument respecting the structure of $\mathcal{M}_{d}(\mathbb{C})$ as a real algebraic variety using Weil restriction. Another difficulty in the length spectrum case is the lack of noetherianity for semialgebraic subsets. We overcome this difficulty by using the notion of admissible subsets introduced in [12].
@@
With the help of Remark 1.3, the proof of Theorem 1.5 in §4 indeed shows that we can weaken the condition of Theorem 1.5 by requiring only that there is a number field $K$ such that for all but finitely many $z_{0}\in{\rm Per}(f)$, we have $\lvert\rho_{f}(z_{0})\rvert^{n(z_{0})}\in K$ for some $n(z_{0})\in\mathbb{Z}_{>0}$.
Finitely many nonzero elements $z_{1},\dots,z_{N}$ in a field $F$ are called multiplicatively independent if for every $N$-tuple $(m_{1},\dots,m_{N})$ of integers, $z_{1}^{m_{1}}\cdots z_{N}^{m_{N}}=1$ if and only if $m_{1}=\cdots=m_{N}=0$. A sequence $(z_{n})_{n=1}^{\infty}$ in $F\setminus\{0\}$ is called multiplicatively independent if all its finite subsequences are multiplicatively independent. Theorem 1.4 immediately implies the existence of infinitely many multipliers for a non-exceptional $f$ whose absolute values are multiplicatively independent.
@@ 1.4.1. Milnor’s conjecture
Milnor [19] showed that an exceptional rational map $f:\mathbb{P}^{1}(\mathbb{C})\to\mathbb{P}^{1}(\mathbb{C})$ of degree $d\geq 2$ must have all its multipliers of periodic points in the ring of integers $\mathcal{O}_{K}$ for some imaginary quadratic number field $K$, and in fact in $\mathbb{Z}$ when $f$ is not a rigid Lattès map. Milnor conjectured that the converse is also true. Milnor’s conjecture was recently proved by Ji-Xie:
@@
See also [5] for a different proof. Recently, Huguin generalized the above result using a different approach:
@@ 1.4.2. A question of Levy and Tucker
On the other hand, at the 2014 AIM workshop Postcritically Finite Maps In Complex And Arithmetic Dynamics, Levy [13] and Tucker [26] independently asked the following question:
@@
Let $f:\mathbb{P}^{1}(\mathbb{C})\to\mathbb{P}^{1}(\mathbb{C})$ be a non-exceptional rational map of degree $d\geq 2$, and let $S$ be the set of all multipliers of periodic points of $f$. Take the subgroup of $\mathbb{C}^{\ast}$ generated by $S\setminus\{0\}$. Is it true that the rank of this group is infinite?
@@ 1.5. Sketch of the proofs
We have already given the idea of Theorem 1.2. Here we explain the proofs of Theorem 1.4 and Theorem 1.5.
We first give the idea of the proof of Theorem 1.5. We argue by contradiction and suppose that $f$ is not exceptional. The first step is to reduce to the case where $f$ is defined over $\overline{\mathbb{Q}}$, which can be done using our Theorem 1.2. After enlarging $K$, we may assume that both $f$ and $\overline{f}$ are defined over $K$, where $\overline{f}$ is the rational map obtained from $f$ by replacing the coefficients with their complex conjugates. In the second step, we combine an arithmetic equidistribution theorem with a result of Zdunik [35] (see [10]) on the Lyapunov exponent to get a contradiction. This argument is inspired by Huguin’s proof of Theorem 1.9. Contrary to the case of Theorem 1.9, we cannot apply the equidistribution theorem to the one-dimensional dynamical system $f:\mathbb{P}^{1}\to\mathbb{P}^{1}$ directly. Our idea is to consider the two-dimensional system $F:=f\times\overline{f}$ on $\mathbb{P}^{1}\times\mathbb{P}^{1}$ instead. More precisely, applying a result of Zdunik [35], we get a sequence $(x_{n})_{n=1}^{\infty}$ of distinct periodic points of $f$ such that
@@
Consider the endomorphism $F:=f\times\overline{f}$ on $\mathbb{P}^{1}\times\mathbb{P}^{1}$ and $\Gamma:=\overline{\{p_{n}=(x_{n},\overline{x_{n}})\}}^{\rm Zar}\subseteq\mathbb{P}^{1}\times\mathbb{P}^{1}$. Without loss of generality, we may assume that $\Gamma$ is irreducible and that the sequence $(p_{n})_{n=1}^{\infty}$ is generic in $\Gamma$. Since the Dynamical Manin-Mumford problem is solved for such an endomorphism $F$ by [8] and $\Gamma$ contains a Zariski-dense subset $\{p_{n}:n\geq 1\}$ of $F$-periodic points, the subscheme $\Gamma$ itself is $F$-periodic. Hence, after replacing $F$ by a suitable iteration, we may assume that $\Gamma$ is $F$-invariant.
Let $\nu_{n}$ be the discrete probability measure equally supported at the union of Galois orbits of iterates of $p_{n}$ under $F$. Then $\nu_{n}$ converges weakly to the canonical measure $\mu$ on $\Gamma$ with respect to $F$ by an equidistribution-type theorem (Theorem 3.1), which is a reformulation of [32, Theorem 3.1]; see §3 for details. Applying $\nu_{n}\to\mu$ to the continuous test functions $\max\{\log\lvert\det(dF)\rvert,A\}$ ($A\in\mathbb{R}$) and letting $A\to$-$\infty$, we get
$2a\leq\int\log\lvert\det(dF)\rvert d\mu,$
which is impossible since the right-hand side equals $2\mathcal{L}_{f}<2a$ by a direct computation.
Next, we sketch the proof of Theorem 1.4. According to [6], postcritically finite (PCF) maps are defined over $\overline{\mathbb{Q}}$ in the moduli space $\mathcal{M}_{d}$ of rational maps of degree $d$, except for the family of flexible Lattès maps. So it suffices to consider the following two cases: 1). $f$ is defined over $\overline{\mathbb{Q}}$, and 2). $f$ is not PCF. For the first case, the conclusion follows from Theorem 1.5. For the second case, we need to develop some new techniques, which are presented in §5. In §5, we consider some pseudo-linear algebra (meaning that the domain of a map may not be the whole vector space) and the vector space $\mathbb{D}(k)_{\mathbb{Q}}=k^{*}\otimes_{\mathbb{Z}}\mathbb{Q}$ for a field $k$ of characteristic zero. We will actually prove a theorem (Theorem 5.7) stronger than the non-PCF case of Theorem 1.4; see §5 and §6 for details. To prove Theorem 5.7, in §6.1 we first deal with the case that $f$ is defined over $\overline{\mathbb{Q}}$. A key ingredient in this step is [3, Lemma 4.1], which is a consequence of Siegel’s theorem on $S$-integral points. The existence of a non-preperiodic critical point of $f$ is essentially used here. In §6.2, we consider the general case and complete the proof. This is achieved by reducing to the case that $f$ is defined over $\overline{\mathbb{Q}}$ via an algebraic-geometric argument and the techniques developed in §5.